OpenStudy (dboyette):

find the area under the standard normal curve for the following: (A) P(z< -0.74) (B) P(-0.87 < z <0) (c) P(-2.03 < z < 1.66)

OpenStudy (zarkon):

just use the standard normal table

OpenStudy (zarkon):

noting that \[\mathbb{P}(a < Z <b)=\mathbb{P}(Z<b)-\mathbb{P}(Z<a)\]

OpenStudy (dboyette):

ok this is where i get confused. Would I look up -0.74? Sorry but I really don't understand this